Find the sec ɵ and tan ɵ if sin ɵ =1/5 and ɵ is in the 1st quadrant
Answer by Alan3354(69443)

Find the sec ɵ and tan ɵ if sin ɵ =1/5 and ɵ is in the 1st quadrant
-------------
cos = sqrt(1 - sin^2) = sqrt(24/25) = sqrt(24)/5
sec = 1/cos = 1/sqrt(24)/5 = 5/sqrt(24)
sec = 5*sqrt(24)/24
sec = 5*sqrt(6)/12
------------
tan = sin/cos = sin*sec
tan = sqrt(6)/12